OpenStudy (anonymous):

Simplify the expressions below & write the final in standard form please :) 1.2x^4(4x^2 + 3x + 1) 2.(4x – 3)(2x^2 – 7x + 1) 3.(x^2 + 4x – 3)(2x^2 + x + 6)

OpenStudy (anonymous):

Help Please..!

OpenStudy (anonymous):

1) = 8x ( 4x^2 + 3x +1)......is this the question?

OpenStudy (anonymous):

or id the term outside the brackets 2x^4 ?

OpenStudy (anonymous):

*is

OpenStudy (anonymous):

Is it 2*4 or 2x*4???

OpenStudy (anonymous):

i'm guessing 2x^4.

OpenStudy (anonymous):

all three of them are questions.

OpenStudy (anonymous):

\[2x ^{4} ( 4x ^{2} + 3x + 1)\] is this the first one?

OpenStudy (anonymous):

@zakiaboo if u don't help us, we can't help u.

OpenStudy (anonymous):

yes @ vaidehi that is the first question.

OpenStudy (anonymous):

then u just need to multiply the 2 terms and use the indices property : \[K ^{A} \times K ^{B} = K ^{A+B}\] I'll do the first one as an example. \[2x ^{4} (4x ^{2} + 3x + 1)\] = \[2x ^{4}(4x ^{2}) + 2x ^{4}(3x) + 2x ^{4}(1)\] \[= 8 x ^{(4+2)} + 6x ^{(4+1)} + 2x ^{4}\] \[= 8x ^{6} + 6x ^{5} + 2x ^{4} + 0x ^{3} + 0x ^{2} + 0x + 0\] this will be the standard form.